# references.bib

### v3

@@ 4.2. Proof of the cohomological integrality theorem
In the rest of this section, we prove Theorem 1.1.
For $\lambda\in X_{*}(T)$, we define the vector space
$\mathcal{H}_{\lambda}\coloneqq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$
By §2.3 We have induction maps $\mathrm{Ind}_{\lambda,\mu}\colon\mathcal{H}_{\lambda}\rightarrow\mathcal{H}_{\mu}$ for $\lambda,\mu\in X_{*}(T)$.
For $\lambda\in X_{*}(T)$, we choose a splitting $\mathfrak{g}^{\lambda}=\mathfrak{g}_{\lambda}\oplus(\mathfrak{g}^{\lambda}/\mathfrak{g}_{\lambda})$. It induces a splitting of the maximal torus $\mathfrak{t}=\mathfrak{g}_{\lambda}\oplus(\mathfrak{t}/\mathfrak{g}_{\lambda})$ (Lemma 3.2). We let $\mathcal{H}_{\lambda}^{\mathrm{prim}}\coloneqq\mathrm{H}^{*}(V^{\lambda}/(G^{\lambda}/G_{\lambda}))$, so that $\mathcal{H}_{\lambda}\cong\mathcal{H}_{\lambda}^{\mathrm{prim}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ (Proposition 3.7). We let $\mathcal{A}_{\lambda}\coloneqq\mathrm{H}^{*}(V^{\lambda}/T)$ and $\mathcal{A}_{\lambda}^{\mathrm{prim}}\coloneqq\mathrm{H}^{*}(V^{\lambda}/(T/G_{\lambda}))$ so that $\mathcal{A}_{\lambda}\cong\mathcal{A}_{\lambda}^{\mathrm{prim}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, $\mathcal{H}_{\lambda}=\mathcal{A}_{\lambda}^{W^{\lambda}}$ and $\mathcal{H}_{\lambda}^{\mathrm{prim}}\cong(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{W^{\lambda}}$.
For each $\lambda\in X_{*}(T)$, we let $J_{\lambda}$ be the smallest $W^{\lambda}$-stable $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of $\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\coloneqq\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$ containing
$k_{\mu,\lambda}=\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V^{\lambda})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g}^{\lambda})}\alpha}$
for all $\mu\in X_{*}(T)$ such that $\lambda\not\preceq\mu$.
We let $\mathcal{A}_{\lambda}^{\mathrm{loc}}\coloneqq\mathcal{A}_{\lambda}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]\cong\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ and $\mathcal{H}_{\lambda}^{\mathrm{loc}}\coloneqq(\mathcal{A}_{\lambda}^{\mathrm{loc}})^{W^{\lambda}}\cong\mathcal{H}_{\lambda}^{\mathrm{prim}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$.
@@
We can characterise $J_{\lambda}$ as the smallest $W^{\lambda}$-stable $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of $\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$ containing
$k_{\mu,\lambda}=\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V^{\lambda})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g}^{\lambda})}\alpha}$
@@
If $\mu\prec\lambda$, then we have $\dim V^{\mu}\leqslant\dim V^{\lambda}$ and $\dim\mathfrak{g}^{\mu}\leqslant\dim\mathfrak{g}^{\lambda}$ with at least one strict inequality. If we have $\lambda\preceq\mu$ then, we have
$\dim V^{\lambda}\leqslant\dim V^{\mu},\quad\dim\mathfrak{g}^{\lambda}\leqslant\dim\mathfrak{g}^{\mu}$
providing a contradiction. Therefore, if $\mu\prec\lambda$, then $\lambda\not\preceq\mu$ and by definition, $k_{\mu,\lambda}\in J_{\lambda}$.
@@
We have $J_{\lambda}^{W^{\lambda}}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$.
@@
$\sum_{\overline{w^{\prime}}\in W^{\lambda}/W^{\mu}}w^{\prime}\cdot\left(\sum_{w\in W^{\mu}}(w\cdot f)\right)k_{\mu,\lambda}$
since $k_{\mu,\lambda}$ is $W^{\mu}$-invariant (Lemma 2.8). This is the formula for
@@
Since $\sum_{w\in W^{\mu}}(w\cdot f)$ is polynomial, its induction $I$ is also polynomial. Therefore, it belongs to $\mathcal{H}_{\lambda}$. Since $I\in J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$ by $W^{\lambda}$-invariance of $J_{\lambda}$, $I\in\mathcal{H}_{\lambda}\cap\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]=\mathcal{H}_{\lambda}^{\mathrm{prim}}$. ∎
@@
Let $w\in W_{\lambda}$. We let $J_{\lambda}^{w}\coloneqq w\cdot J_{\lambda}$. Then, $J_{\lambda}^{w}$ is $W^{\lambda}$-stable by Lemma 4.3. It is an $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of $\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}}(w\cdot\alpha)^{-1}]=\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}}\alpha^{-1}]$. Moreover, it contains
@@
for $\lambda\not\preceq\mu$. Now, we use that by Lemma 2.13, for $w\in W_{\lambda}$, {±$\{\pm k_{\mu,\lambda}\colon\lambda\not\preceq\mu\}=\{\pm k_{w\cdot\mu,w\cdot\lambda}\mid\lambda\not\preceq\mu\}$ since for $w\in W_{\lambda}$, $\overline{w\cdot\lambda}=\overline{\lambda}$. By minimality of $J_{\lambda}$, we deduce that $J_{\lambda}\subseteq J_{\lambda}^{w}$. By symmetry of the argument, we conclude $J_{\lambda}^{w}=J_{\lambda}$. ∎
@@
This follows directly from Lemmas 4.3 and 4.13. ∎
We let $\mathcal{P}_{\lambda}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$ be a direct sum complement of $J_{\lambda}^{W^{\lambda}}$:
@@
$\bigoplus_{\lambda\not\preceq\mu}\mathcal{H}_{\mu}\rightarrow\mathcal{H}_{\lambda}$
@@
This is a calculation very similar to that in the proof of Lemma 4.13. Namely, if $f\in\mathcal{H}_{\mu}$ and $g\in\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, $\mathrm{Ind}_{\mu}(fg)=\mathrm{Ind}_{\mu}(f)g$ by $W^{\lambda}$-invariance of $g$ (Lemma 4.5). Moreover, $fg\in\mathcal{H}_{\mu}$. Therefore, a spanning subset of $J_{\lambda}^{W^{\lambda}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ can be obtained in this way for $\lambda\not\preceq\mu$. Conversely, if $f\in\mathcal{H}_{\mu}\subseteq\mathcal{A}_{\mu}=\mathcal{A}_{\lambda}$ for $\lambda\not\preceq\mu$, we write $f=\sum_{i=1}^{N}f_{i}\otimes g_{i}$ with $f_{i}\in\mathcal{A}_{\lambda}^{\mathrm{prim}}$ and $g_{i}\in\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ so that
$\mathrm{Ind}_{\mu}(f)=\sum_{i=1}^{N}\mathrm{Ind}_{\mu,\lambda}(f_{i})\otimes g_{i}$
@@
$\bigoplus_{\mu\preceq\lambda}\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu})\xrightarrow{\bigoplus_{\mu\prec\lambda}\mathrm{Ind}_{\mu,\lambda}}\mathcal{H}_{\lambda}$
@@
We proceed by induction on $\lambda\in X_{*}(T)$. If $\lambda$ is minimal, then $J_{\lambda}=\{0\}$ and so $\mathcal{P}_{\lambda}=\mathrm{H}^{*}(V^{\lambda}/(G^{\lambda}/G_{\lambda}))$. By Proposition 3.7, the map
@@
Now, let $\lambda\in X_{*}(T)$. We assume that the map of the lemma is surjective for any $\lambda^{\prime}\preceq\lambda$. By associativity of the induction (Proposition 2.7), the image of the map of the lemma coincides with the image of the composition
$\left(\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\prec\mu}\mathcal{P}_{\mu^{\prime}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu^{\prime}})\right)\oplus\left(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\right)\xrightarrow{I\oplus\mathrm{id}}\left(\bigoplus_{\mu\prec\lambda}\mathcal{H}_{\mu}\right)\oplus\left(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\right)\rightarrow\mathcal{H}_{\lambda}$
where the morphism $I$ is $\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\prec\mu}\mathrm{Ind}_{\mu^{\prime},\mu}$. By induction hypothesis,
$\left(\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\prec\mu}\mathcal{P}_{\mu^{\prime}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu^{\prime}})\right)\rightarrow\left(\bigoplus_{\mu\prec\lambda}\mathcal{H}_{\mu}\right)$
@@
For $y\in\mathfrak{t}$, we let $\mathfrak{g}^{y}=\{\xi\in\mathfrak{g}\mid[y,\xi]=0\}$ and $V^{y}=\{v\in V\mid y\cdot v=0\}$.
@@
Via the natural embeddings $X_{*}(T)\subseteq\mathfrak{t}_{\mathbf{Q}}$ and $X^{*}(T)\subseteq\mathfrak{t}_{\mathbf{Q}}^{*}$, the natural pairing $X_{*}(T)\times X^{*}(T)\rightarrow\mathbf{Z}$ corresponds to the natural pairing $\mathfrak{t}\times\mathfrak{t}^{*}\rightarrow\mathbf{Q}$. We let
@@
$\mathcal{W}_{2}\coloneqq\{\alpha\in\mathcal{W}^{\prime}(V)\cup\mathcal{W}^{\prime}(\mathfrak{g})\mid\langle y,\alpha\rangle\neq 0\}$
The vector space $\cap_{\alpha\in\mathcal{W}_{1}}\alpha^{\perp}$ is non-empty, as it contains $y$. The complement of the finite union of rational hyperplanes $U_{y}\coloneqq\cap_{\alpha\in\mathcal{W}_{1}}\alpha^{\perp}\smallsetminus\cup_{\alpha\in\mathcal{W}_{2}}\alpha^{\perp}$ is therefore also non-empty. We let $y^{\prime}\in U_{y}$. For some $N\geqslant 1$, $Ny^{\prime}\in X_{*}(T)\subseteq\mathfrak{t}$. We let $\lambda\coloneqq Ny^{\prime}$. It satisfies the property of the lemma. ∎
@@
By possibly replacing $V$ by $V^{\lambda}$ and $G$ by $G^{\lambda}$, we may assume that $\lambda=0$. It suffices to prove that $J_{\lambda}^{W^{\lambda}}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$ has finite codimension. If we replace $V$ by $V\times\mathfrak{g}$, then the corresponding module $J_{\lambda}\subseteq\mathcal{A}^{\mathrm{prim}}_{\lambda}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}}\alpha^{-1}]$ is smaller than the one for $V$, and therefore so is the corresponding module of invariants $J_{\lambda}^{W^{\lambda}}$, and therefore its codimension is bigger. Moreover, in this case, $J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}$, i.e. there is no need to localise.

### v6

@@ 4.2. Proof of the cohomological integrality theorem
We let $V$ be a weakly symmetric representation of a reductive group $G$. In the rest of this section, we prove Theorem 1.1.
For $\lambda\in X_{*}(T)$, we define the $\mathbf{Z}$-graded vector space
$\mathcal{H}_{\lambda}^{*}\coloneqq\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})\,.$
When $\lambda=0$ is the trivial cocharacter, we write $\mathcal{H}\coloneqq\mathcal{H}_{0}=\mathrm{H}^{*}(V/G)$. More generally, for any pair $\lambda,\mu$, we define $\mathcal{H}_{\lambda,\mu}\coloneqq\mathrm{H}^{*}(V^{\lambda}\cap V^{\mu}/G^{\lambda}\cap G^{\mu})$. By §2.3, we have induction maps $\mathrm{Ind}_{\lambda}\colon\mathcal{H}_{\lambda}\rightarrow\mathcal{H}$ for $\lambda\in X_{*}(T)$ and more generally, $\mathrm{Ind}_{\mu,\lambda}\colon\mathcal{H}_{\lambda,\mu}\rightarrow\mathcal{H}_{\lambda}$ for $\lambda,\mu\in X_{*}(T)$.
For $\lambda\in X_{*}(T)$, we choose a $W_{\lambda}$-invariant splitting $\mathfrak{g}^{\lambda}=\mathfrak{g}_{\lambda}\oplus(\mathfrak{g}^{\lambda}/\mathfrak{g}_{\lambda})$, where $\mathfrak{g}_{\lambda}\coloneqq\mathrm{Lie}(G_{\lambda})$ is the Lie algebra of $G_{\lambda}$4.1). It induces a splitting of the maximal torus $\mathfrak{t}=\mathfrak{g}_{\lambda}\oplus(\mathfrak{t}/\mathfrak{g}_{\lambda})$ (Lemma 3.2). We let
$\mathcal{H}_{\lambda}^{\mathrm{prim}}\coloneqq\mathrm{H}^{*}(V^{\lambda}/(G^{\lambda}/G_{\lambda}^{\circ}))\,$,
so that $\mathcal{H}_{\lambda}\cong\mathcal{H}_{\lambda}^{\mathrm{prim}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ (Proposition 3.7). We let
$\mathcal{A}_{\lambda}\coloneqq\mathrm{H}^{*}(V^{\lambda}/T)\quad\text{and}\quad\mathcal{A}_{\lambda}^{\mathrm{prim}}\coloneqq\mathrm{H}^{*}(V^{\lambda}/(T/G_{\lambda}^{\circ}))$
so that $\mathcal{A}_{\lambda}\cong\mathcal{A}_{\lambda}^{\mathrm{prim}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, $\mathcal{H}_{\lambda}=\mathcal{A}_{\lambda}^{W^{\lambda}}$ and $\mathcal{H}_{\lambda}^{\mathrm{prim}}\cong(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{W^{\lambda}}$.
For each $\lambda\in X_{*}(T)$, we let $J_{\lambda}$ be the smallest $W^{\lambda}$-stable $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of the localisation
$\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\coloneqq\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$
containing
$k_{\mu,\lambda}=\frac{\prod_{\alpha\in\mathcal{W}^{\mu<0}(V^{\lambda})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g}^{\lambda})}\alpha}$
for all $\mu\in X_{*}(T)$ such that $\lambda\not\preceq\mu$ for the order relation $\preceq$ defined in §2.6.
We let
$\mathcal{A}_{\lambda}^{\mathrm{loc}}\coloneqq\mathcal{A}_{\lambda}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]\cong\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$
and
$\quad\mathcal{H}_{\lambda}^{\mathrm{loc}}\coloneqq(\mathcal{A}_{\lambda}^{\mathrm{loc}})^{W^{\lambda}}\cong\mathcal{H}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$
(for this last isomorphism, we use the fact that $W^{\lambda}$ acts trivially on $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ as $G_{\lambda}$ is in the center of $G^{\lambda}$, Lemma 4.4).
@@
We can characterise $J_{\lambda}$ as the smallest $W^{\lambda}$-stable $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of $\mathcal{A}^{\mathrm{prim},\mathrm{loc}}$ containing
$k_{\mu,\lambda}=\frac{\prod_{\alpha\in\mathcal{W}^{\mu<0}(V^{\lambda})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu<0}(\mathfrak{g}^{\lambda})}\alpha}$
@@
If $\mu\prec\lambda$, then we have $\dim V^{\mu}\leqslant\dim V^{\lambda}$ and $\dim\mathfrak{g}^{\mu}\leqslant\dim\mathfrak{g}^{\lambda}$ with at least one strict inequality (where the order relation $\preceq$ is defined in §2.6). If by contradiction we have $\lambda\preceq\mu$ then, we have
$\dim V^{\lambda}\leqslant\dim V^{\mu}\,,\quad\dim\mathfrak{g}^{\lambda}\leqslant\dim\mathfrak{g}^{\mu}\,.$
Therefore, if $\mu\prec\lambda$, then $\lambda\not\preceq\mu$ and by definition, $k_{\mu,\lambda}\in J_{\lambda}$.
@@
We have $J_{\lambda}^{W^{\lambda}}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$. Moreover, $J_{\lambda}\cap\mathcal{H}_{\lambda}^{\mathrm{prim}}=J_{\lambda}^{W^{\lambda}}$.
@@
$\sum_{\overline{w^{\prime}}\in W^{\lambda}/W^{\mu}}w^{\prime}\cdot\left[\left(\sum_{w\in W^{\mu}}(w\cdot f)\right)k_{\mu,\lambda}\right]$
since $k_{\mu,\lambda}$ is $W^{\mu}$-invariant (Lemma 2.10). This is (Proposition 2.11) the formula for
@@
Since $\sum_{w\in W^{\mu}}(w\cdot f)$ is polynomial, its induction $I$ is also polynomial. Therefore, it belongs to $\mathcal{H}_{\lambda}$. Moreover, $I\in J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$ by $W^{\lambda}$-invariance of $J_{\lambda}$. Therefore, $I\in\mathcal{H}_{\lambda}\cap\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]=\mathcal{H}_{\lambda}^{\mathrm{prim}}$.
For the second claim of the lemma, we have obviously $J_{\lambda}^{W^{\lambda}}\subseteq J_{\lambda}$. Combined with the first part of the lemma, this gives the reverse inclusion $J_{\lambda}^{W^{\lambda}}\subseteq J_{\lambda}\cap\mathcal{H}_{\lambda}^{\mathrm{prim}}$. For the direct inclusion, if $f\in J_{\lambda}$ is also in $\mathcal{H}_{\lambda}^{\mathrm{prim}}=(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{W^{\lambda}}$, it clearly is $W^{\lambda}$-invariant: $f\in J_{\lambda}^{W^{\lambda}}$. ∎
@@
Let $w\in W_{\lambda}$. We let $J_{\lambda}^{w}\coloneqq w\cdot J_{\lambda}$. Then, $J_{\lambda}^{w}$ is $W^{\lambda}$-stable by Lemma 4.2. It is an $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of $\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}(w\cdot\alpha)^{-1}]=\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$. Moreover, it contains
@@
for $\lambda\not\preceq\mu$. Now, we use that by Lemma 2.15, for $w\in W_{\lambda}$, $\{\pm k_{\mu,\lambda}\colon\lambda\not\preceq\mu\}=\{\pm k_{w\cdot\mu,w\cdot\lambda}\mid\lambda\not\preceq\mu\}$ since for $w\in W_{\lambda}$, $\overline{w\cdot\lambda}=\overline{\lambda}$. By minimality of $J_{\lambda}$, we deduce that $J_{\lambda}\subseteq J_{\lambda}^{w}$. By symmetry of the argument, we conclude $J_{\lambda}^{w}=J_{\lambda}$. ∎
@@
This follows directly from Lemmas 4.2 and 4.11. ∎
We let $\mathcal{P}_{\lambda}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$ be a $W_{\lambda}$-stable direct sum complement of $J_{\lambda}^{W^{\lambda}}$ (using Corollary 4.12):
@@
$\bigoplus_{\lambda\not\preceq\mu}\mathcal{H}_{\mu,\lambda}\xrightarrow{\bigoplus\mathrm{Ind}_{\mu,\lambda}}\mathcal{H}_{\lambda}$
Similarly, the image of
$\bigoplus_{\mu\prec\lambda}\mathcal{H}_{\mu,\lambda}\xrightarrow{\bigoplus\mathrm{Ind}_{\mu,\lambda}}\mathcal{H}_{\lambda}$
@@
We only prove the second part of the lemma, as the second part can be proved in the exact same way by using Lemma 4.9.
This is a calculation very similar to that in the proof of Lemma 4.11. Namely, if $f\in\mathcal{H}_{\mu,\lambda}$ and $g\in\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, $\mathrm{Ind}_{\mu,\lambda}(fg)=\mathrm{Ind}_{\mu,\lambda}(f)g$ by $W^{\lambda}$-invariance of $g$ (Lemma 4.4). Moreover, $fg\in\mathcal{H}_{\mu,\lambda}$. Therefore, a spanning subset of $J_{\lambda}^{W^{\lambda}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ can be obtained in this way by considering all $\mu\prec\lambda$. Conversely, if $f\in\mathcal{H}_{\mu,\lambda}\subseteq\mathcal{A}_{\mu,\lambda}=\mathcal{A}_{\lambda}=\mathrm{H}^{*}(\mathrm{pt}/T)$ for $\mu\prec\lambda$, we write $f=\sum_{i=1}^{N}f_{i}\otimes g_{i}$ with $f_{i}\in\mathcal{A}_{\lambda}^{\mathrm{prim}}$ and $g_{i}\in\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ so that
$\mathrm{Ind}_{\mu,\lambda}(f)=\sum_{i=1}^{N}\mathrm{Ind}_{\mu,\lambda}(f_{i})\otimes g_{i}$
@@
$\bigoplus_{\begin{subarray}{c}\mu\in X_{*}(T)\\ \mu\preceq\lambda\end{subarray}}\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu})\xrightarrow{\bigoplus_{\mu\preceq\lambda}\mathrm{Ind}_{\mu,\lambda}}\mathcal{H}_{\lambda}$
@@
We proceed by induction on $\lambda\in X_{*}(T)$. If $\lambda$ is minimal, then $J_{\lambda}=\{0\}$ (by Lemma 4.9) and so $\mathcal{P}_{\lambda}=\mathrm{H}^{*}(V^{\lambda}/(G^{\lambda}/G_{\lambda}))$. By Proposition 3.7, the map
@@
Now, let $\lambda\in X_{*}(T)$. We assume that the map of the lemma is surjective for any $\lambda^{\prime}\prec\lambda$. By associativity of the induction (Proposition 2.9), the image of the map of the lemma coincides with the image of the composition
$\left(\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\preceq\mu}\mathcal{P}_{\mu^{\prime}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu^{\prime}})\right)\oplus\left(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\right)\xrightarrow{I\oplus\mathrm{id}}\left(\bigoplus_{\mu\prec\lambda}\mathcal{H}_{\mu}\right)\oplus\left(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\right)\rightarrow\mathcal{H}_{\lambda}$
where the morphism $I$ is $\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\preceq\mu}\mathrm{Ind}_{\mu^{\prime},\mu}$. By induction hypothesis,
$\left(\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\preceq\mu}\mathcal{P}_{\mu^{\prime}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu^{\prime}})\right)\rightarrow\left(\bigoplus_{\mu\prec\lambda}\mathcal{H}_{\mu}\right)$
@@
For $y\in\mathfrak{t}$, we let $\mathfrak{g}^{y}=\{\xi\in\mathfrak{g}\mid[y,\xi]=0\}$ be the centraliser of $y$ in $\mathfrak{g}$ and $V^{y}=\{v\in V\mid y\cdot v=0\}$ where $y$ acts on $V$ via the infinitesimal action.
@@
We may assume that $y\neq 0$. Via the natural embeddings $X_{*}(T)\subseteq\mathfrak{t}_{\mathbf{Q}}$ and $X^{*}(T)\subseteq\mathfrak{t}_{\mathbf{Q}}^{*}$, the natural pairing $X_{*}(T)\times X^{*}(T)\rightarrow\mathbf{Z}$ corresponds to the natural pairing $\langle-,-\rangle\colon\mathfrak{t}\times\mathfrak{t}^{*}\rightarrow\mathbf{C}$ (see also §1.4). We let
@@
$\mathcal{W}_{2}\coloneqq\{\alpha\in\mathcal{W}^{\prime}(V)\cup\mathcal{W}^{\prime}(\mathfrak{g})\mid\langle y,\alpha\rangle\neq 0\}\,.$
The rational vector space $\cap_{\alpha\in\mathcal{W}_{1}}\alpha^{\perp}\subseteq\mathfrak{t}_{\mathbf{Q}}$ is non-zero, as it is defined by the rational equations $\langle-,\alpha\rangle=0$ and its complexification contains $y$. The complement of the finite union of rational hyperplanes $U_{y}\coloneqq\cap_{\alpha\in\mathcal{W}_{1}}\alpha^{\perp}\smallsetminus\cup_{\alpha\in\mathcal{W}_{2}}\alpha^{\perp}$ is therefore also non-empty. We let $y^{\prime}\in U_{y}$ be a rational point ($y\in\mathfrak{t}_{\mathbf{Q}}$). For some $N\geqslant 1$, $Ny^{\prime}\in X_{*}(T)\subseteq\mathfrak{t}$. We let $\lambda\coloneqq Ny^{\prime}$. It satisfies the property of the lemma. ∎
@@
By possibly replacing $V$ by $V^{\lambda}$ and $G$ by $G^{\lambda}$, we may assume that $\lambda=0$. It suffices to prove that $J_{\lambda}^{W^{\lambda}}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$ has finite codimension. If we replace $V$ by $V\times\mathfrak{g}$, then the corresponding module $J_{\lambda}\subseteq\mathcal{A}^{\mathrm{prim}}_{\lambda}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}}\alpha^{-1}]$ is smaller than the one for $V$ (by the very definition of $J_{\lambda}$ at the beginning of §4.2, since when replacing $V$ by $V\times\mathfrak{g}$, the new $k_{\lambda}$ is equal to the numerator of the old $k_{\lambda}$), and therefore so is the corresponding module of invariants $J_{\lambda}^{W^{\lambda}}$, and therefore its codimension is bigger. Moreover, in this case, $J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}$, i.e. there is no need to localise.